The area of a rectangle is 45 square cm. If the length is 4 cm greater than the width, what are the dimensions of the rectangle? I need the algebraic formula, plus the answer.
Answer by mananth(12270)

width x
length = x+4
x(x+4)=45
x^2+4x=45
x^2+4x-45=0
x^2+9x-5x-45=0
x(x+9)-5(x+9)=0
(x+9)(x-5)=0
x= 5 cm the width
length = 5+4 = 9 cm